OpenStudy (anonymous):

Suppose that, from measurements in a microscope, you determine that a certain bacterium covers an area of 1.50 μm2. Convert this to square meters.

OpenStudy (anonymous):

Welcome to OpenStudy! Message me if you have any questions OpenStudy related!

OpenStudy (anonymous):

Convert 5.23×10−6kg/mm3 to kg/m3

OpenStudy (blurbendy):

To your original question Convert \(1.50 \mu m^2\) to \(m^2\) I do unit conversion by fractions. Set up an identity relating the two units: \(1 \mu m= 10^{-6} m\) Divide both sides by the unit you want to get rid of: \(\dfrac{10^{-6}m}{1 \mu m} = 1\) Now we can use that fraction to convert any micrometers we see into meters by multiplying. The value of the fraction is 1, so we don't affect the equation's truth at all, just the units. As we are converting square micrometers to square meters, we will have to multiply by the conversion fraction twice. The units will cancel properly if we set up the equation correctly, and there won't be any question about whether we divided when we should have multiplied, or vice versa. \[1.50 \mu m^2 * \frac{10^{-6}m} {1 \mu m}*\frac{10^{-6}m} {1 \mu m} \]Observe that the units cancel the way we want:\[1.50 \cancel{\mu m^2} * \frac{10^{-6}m} {1 \cancel{\mu m}}*\frac{10^{-6}m} {1 \cancel{\mu m}} = 1.50 * 10^{-6} * 10^{-6} m^2\] Stolen from my friend @whpalmer4